# arXiv:2602.14013v2[math.CV] 01 Mar 2026

### v2

@@ 3 On image of the continuum under open discrete and closed mappings
Since $P$ in (3.5) may be done arbitrary big, the relations (3.5) and (3.11) contradict each other. This completes the proof for the case when the relation (1.6) holds. Observe that, the case when $Q$ has a finite mean oscillation in $A$ is a particular case of the relation (1.6) (see the last part of the proof of Theorem 5.5 in [Sev2]). The proof is wholly complete. $\Box$
Corollary 3.1.   Assume that, under conditions of Lemma 3, $A_{0}$ is some another compactum in $D_{0}.$ Then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
Proof.   Since $D_{m}$ converges to $D_{0}$ as its kernel, there is $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}.$ It remains to show that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}$ and some $\delta_{2}>0.$ Assume the contrary, i.e., $h(f_{m}(A_{0}),\partial f_{m}(D_{m})\rightarrow 0$ as $m\rightarrow\infty.$ Now, $h(f_{m}(A_{0}),\partial f_{m}(D_{m})=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty$ for $x_{m}\in A_{0}$ and $y_{m}\in f_{m}(D_{m}).$ We may consider that $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ $x_{0}\in A_{0}.$ Put $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}.$ We join $x_{0}$ with $A$ by a path $\gamma$ in $D_{0}.$ Now, $B:=A\cup|\gamma|\cup\overline{B(x_{0},\varepsilon_{0})}$ is a continuum in $D_{0}$ such that $B\subset D_{m}$ for sufficiently large $m,$ $x_{m}\in B$ for sufficiently large $m$ and $h(f_{m}(B))\geqslant h(f_{m}(B))\delta$ for above $m.$ By Lemma 3, $h(f_{m}(B),\partial f_{m}(D_{m}))\geqslant\delta_{*}>0$ for $m\geqslant m_{3},$ $m_{3}\in{\mathbb{N}}.$ However, the latter contradicts with $h(f_{m}(A_{0}),\partial f_{m}(D_{m})=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty,$ because obviously $h(f_{m}(B),\partial f_{m}(D_{m}))\leqslant h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ The obtained contradiction completes the proof. $\Box$
4 Proof of the main results
Proof of Theorem 1. Let us prove the theorem by contradiction. Assume the assertion of Theorem 1 is false. Then for some $\varepsilon_{0}>0$ and for numbers $\delta_{k}=\frac{1}{k},$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},y_{k}\in D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(y_{k}))\geqslant\varepsilon_{0}$ whenever $|x_{k}-y_{k}|<\frac{1}{k}.$ By Lemma 2 every $f_{m_{k}}$ has a continuous extension to $\partial D_{m_{k}},$ so that $f_{m_{k}}$ is uniformly continuous in $\overline{D_{m_{k}}}.$ Thus, we may consider that the sequence $m_{k},$ $k=1,2,\ldots,$ is increasing. Without loss of generality, we may consider that the above holds for $m=1,2,\ldots,$ but not for $m_{k}.$ In particular,
@@ 4 Proof of the main results
Now, by the triangle inequality,
@@
By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ There are two cases: 1) $x_{0}\in D_{0},$ 2) $x_{0}\in\partial D_{0}.$ In the first case, when $x_{0}$ is the inner point of $D_{0},$ all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS, Theorems 6.1, 6.5] for $p=n$ and [Sev2, Theorems 4.1, 4.3] for n1<p<n.n-1<p<n. In this case, the right side in (4.2) tends to zero as $m\rightarrow\infty$ which contradicts with (4.1), and the proof finishes.
@@
Let us consider the case 2) $x_{0}\in\partial D_{0}.$ Observe that $A_{m}:=A\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0})$ for some $\varepsilon_{0}>0,$ because the $A$ is a continuum in $D_{0}.$ Thus, all of conditions of Theorem 2 are satisfied. In particular, it follows from this theorem that $f_{m}$ has a continuous extension to $\partial D_{m}$ for any $m\in{\mathbb{N}}.$ Thus, there is $x_{m}\in D_{m}$ such that $h(f_{m}(x^{\,\prime}_{m}),f_{m}(x_{0}))<\frac{1}{m}.$ Now, by the triangle inequality, by (4.2) and by Theorem 2 we obtain that
$h(f_{m}(x_{m}),f_{m}(y_{m}))\leqslant h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(x_{0}))+$
$+h(f_{m}(x_{0}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(y_{m}))<$ (4.3)
$+h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(y_{m}))+\frac{2}{m}\rightarrow 0\qquad\text{as}\qquad m\rightarrow\infty.$
The relations (4.1) and (4.3) contradict each other and prove the uniformly equicontinuity of $f_{m}.$
@@
Given a compactum $K$ in $D_{0},$ there is $m_{0}=m_{0}(K)$ such that $K\subset D_{m}$ for $m\geqslant m_{0}(K).$ Now, by the proving above, $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous family on $K.$ By Arzela-Ascoli Theorem (see. e.g., [Va, Theorem 20.4]) $f_{m}$ forms a normal family on $K,$ i.e., there exists a subsequence $f_{m_{k}},$ $k=1,2,\ldots,$ such that $f_{m_{k}}(x)\rightarrow f(x)$ as $k\rightarrow\infty.$ Since $K$ is an arbitrary compactum in $D_{0},$ the mapping $f(x)$ is defined in all $D_{0}.$ Now, we prove that $f$ has a continuous extension at every point $x_{0}\in\partial D_{0}.$ By the proving above, given $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$ Fix $x_{0}\in\partial D_{0}.$ Let $\delta_{*}=\delta_{*}(\varepsilon)=\delta/2$ and let $x,y\in B(x_{0},\delta_{*})\cap D_{0}.$ Now, by the triangle inequality $|x-y|<2\delta_{*}=\delta.$ Since $x,y\in D_{0},$ there is $m_{0}=m_{0}(x,y)\in{\mathbb{N}}$ such that $x,y\in D_{0}$ for every $m\geqslant m_{0}.$ Taking into account the mentioned above, we obtain that $h(f_{m}(x),f_{m}(y))<\varepsilon,$ $m\geqslant m_{0}.$ Taking the limit as $m\rightarrow\infty$ in the latter inequality, we obtain that $h(f(x),f(y))\leqslant\varepsilon.$ Now, $f$ satisfies the Cauchy condition at $x_{0},$ so that $f$ has a limit at $x_{0},$ as required.
The existing of $\delta_{1}>0$ for which $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ follows by Lemma 3. The existing of $\delta_{2}>0$ for which $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for any $m=1,2,\ldots$ follows by Corollary 3.  $\Box$
@@ 5 Main Lemma for mappings with a branching
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ the family of all open discrete closed ring mappings $f:D_{f}\rightarrow\overline{{\mathbb{R}}^{n}}\setminus E$ defined in some a domain $D_{f}\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ such that the following condition holds: for any domain $D^{\,\prime}_{f}:=f(D_{f})$ there exists a continuum $K_{f}\subset D^{\,\prime}_{f}$ such that $h(K_{f})\geqslant\delta$ and $h(f^{\,-1}(K_{f}),\partial D_{f})\geqslant\delta>0.$ The following statement holds.
The following lemma is an analog of Lemma 2 for mappings which are not homeomorphisms.
Lemma 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
@@
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ for $f_{m};$
@@
Denote by $C_{m}$ the image of the path $\gamma_{m}$ under the mapping $f_{m}.$ It follows from relation (2.5) that a condition of the form (2.6) is satisfied, where $h$ denotes the chordal diameter of the set.
By the definition of the family $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D}),$ for any $f_{m}\in\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ and any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Since by the condition of the lemma the domains $D^{\,\prime}_{m}$ form the equi-uniform family with respect to $p$-modulus and due to (2.6), we obtain that the inequality
@@
holds for all $m=1,2,\ldots$ and some $b>0.$ Consider the family $\Gamma_{m},$ consisting of all paths $\beta:[0,1)\rightarrow D^{\,\prime}_{m},$ where $\beta(0)\in C_{m}$ and $\beta(t)\rightarrow p\in K_{m}$ for $t\rightarrow 1.$ Let $\Gamma^{*}_{m}$ be a family of all total $f_{m}$-liftings $\alpha:[0,1)\rightarrow D$ of the family $\Gamma_{m}$ with origin on $\gamma_{m}$ (see Proposition 3). Since $f_{m}$ is closed, we have: $\alpha(t)\rightarrow f^{\,-1}_{m}(K_{m}),$ where $f^{\,-1}_{m}(K_{m})$ is the total preimage of the continuum $K_{m}$ under the mapping $f_{m}.$ By (1.3), it follows that
@@
$\leqslant M_{p}(f_{m}(\Gamma(|\gamma_{m}|,f^{\,-1}_{m}(K_{m}),D_{m})))\leqslant\int\limits_{A(x_{0},\frac{1}{2^{m}},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (5.2)
for each measurable function $\eta:(\frac{1}{2^{m}},\varepsilon_{0})\rightarrow[0,\infty],$ such that $\int\limits_{\frac{1}{2^{m}}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Note that, the function
@@
where $I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt,$ satisfies the normalization condition (1.4) with $r_{1}:=2^{\,-m},$ $r_{2}:=\varepsilon_{0}.$ Therefore from the conditions (2.3)–(2.4) and (5.2) it follows that
@@
However, relations (5.3) and (5.4) contradict with (5.1). The resulting contradiction indicates that the assumption (2.5) was incorrect, and, therefore, the family $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ is equicontinuous at $x_{0}\in\partial D.$ $\Box$
Theorem 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
@@
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ for $f_{m};$

### v3

@@ 3 On image of the continuum under open discrete and closed mappings
Since $P$ in (3.2) may be done arbitrary large, the relations (3.2) and (3.8) contradict each other. This completes the proof for the case when the relation (1.5) holds. Observe that, the case when $Q$ has a finite mean oscillation in $A$ is a particular case of the relation (1.5) (see the last part of the proof of Theorem 5.5 in [Sev1]). The proof is wholly complete. $\Box$
Corollary 3.1.   Assume that, under conditions of Lemma 3, $A_{0}$ is some another compactum in $D_{0}.$ Then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
Proof.   Since $D_{m}$ converges to $D_{0}$ as its kernel, there is $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}.$ It remains to show that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}$ and some $\delta_{2}>0.$ Assume the contrary, i.e., $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\rightarrow 0$ as $m\rightarrow\infty.$ Now, $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty$ for $x_{m}\in A_{0}$ and $y_{m}\in f_{m}(D_{m}).$ We may consider that $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ $x_{0}\in A_{0}.$ Put $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}.$ We join $x_{0}$ with $A$ by a path $\gamma$ in $D_{0}.$ Now, $B:=A\cup|\gamma|\cup\overline{B(x_{0},\varepsilon_{0})}$ is a continuum in $D_{0}$ such that $B\subset D_{m}$ for sufficiently large $m,$ $x_{m}\in B$ for sufficiently large $m$ and $h(f_{m}(B))\geqslant h(f_{m}(B))\geqslant\delta$ for above $m.$ By Lemma 3, $h(f_{m}(B),\partial f_{m}(D_{m}))\geqslant\delta_{*}>0$ for $m\geqslant m_{3},$ $m_{3}\in{\mathbb{N}}.$ However, the latter contradicts with $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty,$ because obviously $h(f_{m}(B),\partial f_{m}(D_{m}))\leqslant h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ The obtained contradiction completes the proof. $\Box$
4 Proof of Theorem 1
The obtained contradiction completes the proof. $\Box$
Proof of Theorem 1. I. Let us prove the item I of Theorem 1 by the contradiction. Assume that the assertion of Theorem 1 is false. Then for some $\varepsilon_{0}>0$ and for numbers $\delta_{k}=\frac{1}{k},$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},y_{k}\in D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(y_{k}))\geqslant\varepsilon_{0}$ whenever $|x_{k}-y_{k}|<\frac{1}{k}.$ By Lemma 2 every $f_{m_{k}}$ has a continuous extension to $\partial D_{m_{k}},$ so that $f_{m_{k}}$ is uniformly continuous in $\overline{D_{m_{k}}}.$ Thus, we may consider that the sequence $m_{k},$ $k=1,2,\ldots,$ is increasing. Without loss of generality, we may consider that the above holds for $m=1,2,\ldots,$ but not for $m_{k}.$ In particular,
@@ 4 Proof of Theorem 1
By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ There are two cases: 1) $x_{0}\in D_{0},$ 2) $x_{0}\in\partial D_{0}.$ In the first case, when $x_{0}$ is the inner point of $D_{0},$ all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS1, Theorems 6.1, 6.5] for $p=n$ and [Sev1, Theorems 4.1, 4.3] for n1<p<n.n-1<p<n. Now, by the triangle inequality,
@@
In this case, the right side in (4.2) tends to zero as $m\rightarrow\infty$ which contradicts with (4.1), and the proof finishes.
@@
Let us consider the case 2) $x_{0}\in\partial D_{0}.$ Observe that $A_{m}:=A\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0})$ for some $\varepsilon_{0}>0,$ because the $A$ is a continuum in $D_{0}.$ Thus, all of conditions of Theorem 2 are satisfied. By this theorem, for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$ In particular, it follows from this theorem that there is $M=M(\varepsilon)\in{\mathbb{N}}$ such that
$h(f_{m}(x_{m}),f_{m}(y_{m}))<\varepsilon\,,\qquad m\geqslant M(\varepsilon)\,.$ (4.3)
The relations (4.1) and (4.3) contradict each other and prove the uniform equicontinuity of $f_{m}.$
@@
II. Given a compactum $K$ in $D_{0},$ there is $m_{0}=m_{0}(K)$ such that $K\subset D_{m}$ for $m\geqslant m_{0}(K).$ Now, by the proving above, $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous family on $K.$ By Arzela-Ascoli Theorem (see. e.g., [Va, Theorem 20.4]) $f_{m}$ forms a normal family on $K,$ i.e., there exists a subsequence $f_{m_{k}},$ $k=1,2,\ldots,$ such that $f_{m_{k}}(x)\rightarrow f(x)$ as $k\rightarrow\infty.$ Since $K$ is an arbitrary compactum in $D_{0},$ the mapping $f(x)$ is defined in all $D_{0}.$ Now, we prove that $f$ has a continuous extension at every point $x_{0}\in\partial D_{0}.$ By the proving above, given $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$ Fix $x_{0}\in\partial D_{0}.$ Let $\delta_{*}=\delta_{*}(\varepsilon)=\delta/2$ and let $x,y\in B(x_{0},\delta_{*})\cap D_{0}.$ Now, by the triangle inequality $|x-y|<2\delta_{*}=\delta.$ Since $x,y\in D_{0},$ there is $m_{0}=m_{0}(x,y)\in{\mathbb{N}}$ such that $x,y\in D_{0}$ for every $m\geqslant m_{0}.$ Taking into account the mentioned above, we obtain that $h(f_{m}(x),f_{m}(y))<\varepsilon,$ $m\geqslant m_{0}.$ Taking the limit as $m\rightarrow\infty$ in the latter inequality, we obtain that $h(f(x),f(y))\leqslant\varepsilon.$ Now, $f$ satisfies the Cauchy condition at $x_{0},$ so that $f$ has a limit at $x_{0},$ as required.
Let $A_{0}$ be a compactum in $D_{0}$ and let $G$ be an arbitrary domain in $D_{0}$ with $A_{0}\cup A\overline{G}\subset D_{0}.$ Since $D_{m}$ converges to $D_{0}$ as its kernel, there is $m_{0}\in{\mathbb{N}}$ such that $G\cup A_{0}\subset D_{m},$ $m\geqslant m_{0}.$ Due to the above, $f_{m_{k}}$ converges to $f$ uniformly in $G.$ Now, $f$ is either a homeomorphism, or a constant in $G$ (see [RS2, Theorems 4.1 and 4.2] for $p=n$ and [Cr2, Theorem 1] for n1<p<nn-1<p<n). Due to the condition $h(f_{m}(A))\geqslant\delta,$ $f$ is homeomorphism in $G.$ Since $G$ is arbitrary compact subdomain in $D_{0},$ $f$ is continuous and injective in $D_{0}.$ Now, $f$ is a homeomorphism in $D_{0},$ as required (see, e.g., [RS2, Corollary 3.1]).
Let us to prove that, there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$ Without loss of generality, we may consider that $f_{m}$ converges to $f$ locally uniformly as $m\rightarrow\infty.$ Let $\varepsilon:=\frac{1}{2}h(f(A_{0}),\partial f(G))$ ($\varepsilon>0$ because $f$ is a homeomorphism in $G$ and, consequently, $f(A_{0})\in\textrm{Int}\,f(G)$). By the triangle inequality and due to the uniform convergence of $f_{m}$ to $f,$ given $\varepsilon>0$ there is $M=M(\varepsilon)$ such that $h(f_{m}(x),f(x))<\varepsilon/2$ for any $x\in A_{0}.$ Let $y\in\partial f(G).$ By the triangle inequality
$h(f_{m}(x),y)\geqslant h(y,f(x))-h(f(x),f_{m}(x))\geqslant\varepsilon-\varepsilon/2=\varepsilon/2=\frac{1}{4}h(f(A_{0}),\partial f(G))\,.$
Thus, taking the $\inf$ over $x\in A_{0}$ and $y\in\partial f(G),$ we obtain that $h(f_{m}(A_{0}),\partial f(G))\geqslant\frac{1}{4}h(f(A_{0}),\partial f(G)):=\delta_{*}.$ Let now
$h(f_{m}(A_{0}),\partial f_{m}(G))=h(f_{m}(A_{0}),f_{m}(\partial G))=h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))\,$,
where $x_{m}\in A_{0}$ and $x^{\,\prime}_{m}\in\partial G.$ Now, by the triangle inequality and by the locally uniform convergence of $f_{m}$ to $f,$ we obtain that
$h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))\geqslant h(f_{m}(x_{m}),f(x^{\,\prime}_{m}))-h(f(x^{\,\prime}_{m}),f_{m}(x^{\,\prime}_{m}))\geqslant\delta_{*}-\delta_{*}/2=\delta_{*}/2>0$
for sufficiently large $m\in{\mathbb{N}}.$ The latter inequality implies that $h(f_{m}(A_{0}),\partial f_{m}(G))\geqslant\delta_{*}/2:=\delta_{2}.$ Finally, since $G\subset D_{m}$ for sufficiently large $m\in{\mathbb{N}},$ by Proposition 4 we obtain that $h(f_{m}(A_{0}),\partial f_{m}(G))\geqslant\delta_{2}$ for above $m,$ as required.
III. Let $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $x_{m}\rightarrow x_{0}$ and $f_{m}\rightarrow f$ locally uniformly in $D_{0}$ as $m\rightarrow\infty.$ We need to show that $f_{m}(x_{m})\rightarrow f(x_{0}).$ If $x_{0}\in D_{0},$ the relation $f_{m}(x_{m})\rightarrow f(x_{0})$ as $m\rightarrow\infty$ directly follows from the locally uniform convergence of $f_{m}$ to $f$ in $D_{0}.$ Now, let $x_{0}\in\partial D_{0}.$ We lead the proof by the contradiction, i.e., assume that $h(f_{m_{k}}(x_{m_{k}}),f(x_{0}))\geqslant\varepsilon_{0}>0$ for some an increasing sequence of numbers $m_{k}=1,2,\ldots,$ and some $\varepsilon_{0}>0.$ Again, going to renumbering, if required, we may assume that
$h(f_{m}(x_{m}),f(x_{0}))\geqslant\varepsilon_{0}>0\,,\qquad m=1,2,\ldots\,.$ (4.4)
Consider some a sequence $y_{m}\rightarrow x_{0},$ $y_{m}\in D_{0},$ $m\rightarrow\infty.$ Now, given $m\in{\mathbb{N}},$ by the locally uniform convergence of $f_{m}$ to $f,$ there is $k=k_{m}$ such that $h(f_{k_{m}}(y_{k_{m}}),f(y_{k_{m}}))<1/m\,.$ We may consider that $k_{m}$ is an increasing sequence by $m=1,2,\ldots.$ Let $z_{m}:=y_{k_{m}}.$ Now, $z_{m}\rightarrow x_{0}$ and $f(z_{m})\rightarrow f(z_{0})$ as $m\rightarrow\infty$ because $f$ is continuous at $x_{0}$ (see item II). Now, by the triangle inequality and I we obtain that
$h(f_{k_{m}}(x_{k_{m}}),f(x_{0}))\leqslant h(f_{k_{m}}(x_{k_{m}}),f_{k_{m}}(z_{m}))$
$+h(f_{k_{m}}(z_{m}),f(z_{m}))+h(f(z_{m}),f(x_{0}))\rightarrow 0$ (4.5)
as $m\rightarrow\infty.$ The relation (4.5) contradicts with (4.4), that proves that $f_{m}(x_{m})\rightarrow f(x_{0})$ as $m\rightarrow\infty.$ Finally, $f(x_{0})\in\partial f(D_{0})$ whenever $x_{0}\in\partial D_{0},$ because $f$ is a homeomorphism in $D_{0}$ (see, e.g., [MRSY, Proposition 13.5]). Theorem is completely proved. $\Box$
@@ 5 Main Lemma for mappings with a branching
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ the family of all open discrete closed ring mappings $f:D_{f}\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some a domain $D_{f}\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ such that the following condition holds: for any domain $D^{\,\prime}_{f}:=f(D_{f})$ there exists a continuum $K_{f}\subset D^{\,\prime}_{f}$ such that $h(K_{f})\geqslant\delta$ and $h(f^{\,-1}(K_{f}),\partial D_{f})\geqslant\delta>0.$ The following lemma is an analog of Lemma 2 for mappings which are not homeomorphisms.
Lemma 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
@@
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ for $f_{m};$
@@
Denote by $C_{m}$ the image of the path $\gamma_{m}$ under the mapping $f_{m}.$ It follows from relation (2.3) that a condition (2.4) is satisfied, where $h(C_{m})$ denotes the chordal diameter of the set $C_{m}.$
By the definition of the family $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D}),$ for any $f_{m}\in\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ and any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Since by the condition of the lemma the domains $D^{\,\prime}_{m}$ form the equi-uniform family with respect to $p$-modulus and due to (2.4), we obtain that the inequality
@@
holds for all $m=1,2,\ldots$ and some $b>0.$ Consider the family $\Gamma_{m},$ consisting of all paths $\beta:[0,1)\rightarrow D^{\,\prime}_{m},$ where $\beta(0)\in C_{m}$ and $\beta(t)\rightarrow p\in K_{m}$ as $t\rightarrow 1.$ Let $\Gamma^{*}_{m}$ be a family of all total $f_{m}$-liftings $\alpha:[0,1)\rightarrow D$ of the family $\Gamma_{m}$ starting at $|\gamma_{m}|$ (see Proposition 3). Since $f_{m}$ is closed, we have: $\alpha(t)\rightarrow f^{\,-1}_{m}(K_{m}),$ where $f^{\,-1}_{m}(K_{m})$ is the total preimage of the continuum $K_{m}$ under the mapping $f_{m}.$ By (1.3), it follows that
@@
$\leqslant M_{p}(f_{m}(\Gamma(|\gamma_{m}|,f^{\,-1}_{m}(K_{m}),D_{m})))\leqslant\int\limits_{A(x_{0},2^{\,-m},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (5.2)
for each measurable function $\eta:(2^{\,-m},\varepsilon_{0})\rightarrow[0,\infty],$ such that $\int\limits_{2^{\,-m}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Note that, the function
@@
where $I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt,$ satisfies the normalization condition (1.4). Therefore from the conditions (2.1)–(2.2) and (5.2) it follows that
@@
However, the relation (5.3) contradicts with (5.1) and (5.4). The resulting contradiction indicates that the assumption (2.3) was incorrect, and, therefore, the family $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ is equicontinuous at $x_{0}\in\partial D.$ $\Box$
Theorem 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
@@
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ for $f_{m};$